# The Failed Migration of Academic Twitter

### v2

@@ Abstract.
Following changes in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some. Our study looks at the dynamics of this migration. Utilizing publicly available user account data, we track the posting activity of academics on Mastodon over a one year period. We also gathered follower-followee relationships to map internal networks, finding that the subset of academics who engaged in migration were well-connected. However, this strong internal connectivity was insufficient to prevent users from returning to Twitter/X. Our analyses reveal significant challenges sustaining user engagement on Mastodon due to its decentralized structure as well as competition from other platforms such as Bluesky and Threads. The movement lost momentum after an initial surge of enthusiasm where the main network was fully established as most users did not maintain their activity levels, and those who did faced lower levels of engagement. Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focus on community building for long-term user engagement.
Social Media, Mastodon, Twitter, Social Movement, Computational Social Science, Decentralized Platforms
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Social networking sitesccs: Human-centered computing Social mediaccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Social networks
1. Introduction
@@
Social media platforms have become essential outlets for communication, information dissemination, and professional networking. Among these platforms, Twitter (now known as “X”) has gained significant popularity and has been likened to a virtual town square for its role facilitating public discourse (Mascaro and Goggins, 2012). Academics and journalists populate sub-communities on Twitter and use the platform for professional communication (Jurkowitz and Gottfried, 2022; Malik et al., 2019). Academics have extensively used Twitter to share research findings, announce academic events such as conferences, and contribute to wider scholarly discourse (Mohammadi et al., 2018; Ebner, 2013; Mahrt et al., 2014). However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives (Cava et al., 2023).
@@
Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their areas of interest. Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network (Zignani et al., 2018). This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control (Raman et al., 2019). Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels (Raman et al., 2019).
@@
The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities. Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014). Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts of change (Fiesler and Dym, 2020).
Additionally, research has shown that user migration is influenced by perceived platform attractiveness and peer effects (Hou and Shiau, 2020). A study of users migrating from Twitter to Mastodon post-Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023). Analyses showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023). A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
@@
Despite prior work characterizing the general population of platform users, there remains a gap in targeted research concerning specific professional communities. In this study, we address this gap by examining the migration patterns and behaviors of academic users, providing a more nuanced understanding of how these particular communities navigate platform transitions. Given the significant role that Twitter plays in an academic’s professional life (Knight and Kaye, 2016; Klar et al., 2020; Stewart, 2015), academics represent a uniquely valuable group for studying the migratory dynamics. Our study addresses the following research questions:
RQ1: To what extent were academics engaged in the migration from Twitter to Mastodon, and how can the success or failure of this migration be assessed?
RQ2: Which academic fields have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
@@
Our findings indicate that majority of academic researchers who moved to Mastodon failed to maintain their posting activity and follower engagement over long term. Although migrated academic accounts formed strongly-connected internal networks, these connections were insufficient to sustain user engagement over time. Of note, however, communities that established field specific server instances had higher user retention rate compared to those on general-purpose servers. Cross platform analysis of most recent Twitter activity showed that most users returned to Twitter. These results underscore the challenges involved and the potential opportunities to build and foster sustainable communities at scale on new platforms.
2. Related Work
2.1. Characteristics of Mastodon
The decentralized structure Mastodon and similar platforms leads to unique user engagement dynamics that are different from centralized platforms. Research suggests that users are drawn to Mastodon for the ease of protecting their information from data mining and the absence of algorithms shaping their feeds (Lee and Wang, 2023). However, this can also make content discovery more difficult (Zignani et al., 2019). The decentralized structure encourages strong, tight-knit communities within individual server instances but fewer connections across instances (Zignani et al., 2018). These characteristics pose challenges for migration as well.
2.2. Cross-platform migration
Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014). The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities’ retention.
@@ 2.2. Cross-platform migration
Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process (Fiesler and Dym, 2020). Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence (Hou and Shiau, 2020). Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023). Analysis showed Twitter to Mastodon migration is characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023). A similar study on Twitter-to-Mastodon migration reveals lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
Our work looks more deeply at the specific case of the relatively well-synchronized and coordinated movement of communities of academic users across platforms.
3. Data
Note on terminology: Twitter rebranded to X in July 2023, which occurred midway through our data collection. For simplicity, we use the terminology Twitter and tweets throughout the paper.
@@ 3. Data
Our analyses are primarily based on a public Github repository containing lists of academics on Mastodon [^1] https://github.com/nathanlesage/academics-on-mastodon. These lists were curated by researchers from 50 distinct disciplines, who voluntarily added their own names and accounts. Each list corresponds to a research discipline, containing account details of researchers within that area, their Mastodon IDs, and a short description of their research. Some lists also included each user’s Twitter handle. While the database continues to expand, for the purpose of longitudinal analysis, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022. That is, the users in our dataset were part of a coordinated, synchronized migration effort in November 2022.
@@
In total, we collected the Mastodon profiles and corresponding public metrics, e.g., number of followers, following, timestamps of most recent status of 7,542 academic user accounts using a customized Mastodon API [^2] https://github.com/halcy/Mastodon.py.git. Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of user activity and engagement over the span of a year. In addition, we collected the most recent Mastodon data related to follower and following accounts for each scholar to conduct network analyses based on fields of study and associated servers.
@@
Finally, to understand discussions surrounding Mastodon in academic context, we collected tweets during the one year window via Zeeschuimer [^3] https://github.com/digitalmethodsinitiative/zeeschuimer using the boolean search “mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01”. Since many tweets were included because the username contained the word “mastodon” (as some users use their Mastodon account name as their Twitter username) we further filtered out these tweets. We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets.
Figure 1. Waterfall plot showing number of monthly active users (active=at least one post per month).
4. Results
@@
Following, we present findings from multiple analyses conducted to understand the dynamics of the migration. First, we used the longitudinal data from Mastodon to understand user attrition and engagement trends over the year. Then, we used network analysis to understand the connectivity and distribution of academics across Mastodon servers. Next, we did a cross-platform analysis using the subset of users who provided Twitter handles comparing their activity patterns on Twitter and Mastodon. Finally, we ran exploratory analyses on Twitter posts related to the migration to understand user attitudes towards migration.
4.1. Longitudinal Analysis
4.1.1. Attrition of academic accounts on Mastodon
@@
Figure 1 gives a count waterfall plot illustrating the change in number of active users aggregated by month and a line plot showing the monthly retention rate based on the previous month. In this context, we considered a user to be active if they made at least a post in a given month. In the initial month of data collection (November 2022), there were 7,505 active scholars who had posted within that same month. [^4] There were 1, 4, and 32 scholars had their last activity in April, May, and October 2022, respectively, which fell outside our collection window. These scholars were excluded from the waterfall plot but included in subsequent analysis. Over the next seven months, the number of active scholars experienced a decline, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023. We observe a subtantial rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
4.1.2. Engagement with academic accounts on Mastodon
@@
Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. To do so, we first characterize accounts based on their activity.
@@
Active users. We define active users as scholars who engaged on Mastodon during the initial Twitter-to-Mastodon migration, which began in late October 2022. For our purposes, we’ve defined this initial period between October 15 and November 30, 2022. We further divide active users into sub-categories based on level of engagement: one-time visitors; short-term, long-term, and persistent adopters. Each category reflects the duration of their activity on Mastodon following the migration.
One-time visitors. These academics created an account on Mastodon at the beginning of the Twitter-to-Mastodon migration, posted only within the first week of joining and then abandoned the platform ($n_{days}\leq 7$). This brief engagement suggests a possible exploration of the platform following the migration advocacy, without finding enough value or community to continue their activity.
Short-term adopters. This group includes those who actively used Mastodon for more than one week but did not persist beyond 100 days. They left the platform possibly due to unmet expectations or a stronger pull back to more familiar platforms like Twitter.
Long-term adopters. These users remained active on Mastodon from the time of migration for relatively long time ($100<n_{days}\leq 300$) but have been inactive.
@@
Persistent adopters. These users remained active on Mastodon from initial migration though the end of our analyses ($n_{days}>300$). Their sustained engagement indicates successful integration into the Mastodon community.
4.1.3. Inactive users.
@@
Academics in this category were already using Mastodon before the migration wave from Twitter. Despite their early presence on the platform, they remained inactive during the migration period. One user was inactive during the first month but engaged briefly afterward. Inactive users are not included in subsequent analyses. The distribution of users is given in Table 1. Note that only 35.4% of academic users in our dataset are active and consistent in their use of Mastodon. The majority demonstrate sporadic engagement.
@@
Active One-time visitors 1,183
Short-term adopters 1,624
Long-term adopters 2,065
Persistent adopters 2,665
@@
Table 1. Number of users, by activity.
@@
We calculate the average following, follower and post counts for each user category, aggregated by month show in Figure 2. Our analysis reveals a sharp increase in all three public metrics at the onset of the migration period, specifically from November 2022 to December 2022. However, beyond this period, the rate of increase for following and follower metrics slows considerably. For one-time, short-term, and long-term users, growth in following and follower counts stagnated beginning in early 2023. As these users were not regularly contributing content, their visibility and potential to attract followers were likewise limited. In contrast, persistent users continued to post regularly. Despite their consistent activity, however, growth in their connections remained slow. This slow growth highlights the challenges faced in establishing and expanding social ties and communities on Mastodon. The persistent users’ continued efforts to engage and post suggest a commitment to the platform, yet the limited increase in followers indicates potential barriers in user retention and community building.
Figure 2. Average public metrics and statuses count for active users.
@@
4.1.4. Field-specific activity differences among academic accounts on Mastodon
@@
As discussed in Section 3, users self-identified their research disciplines in public lists. We aggregate these disciplines into 9 broad fields, consistent with existing taxonomies of academic disciplines (bepress., 2019). Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers. Inactive users were excluded from subsequent analyses. Disciplinary statistics of the 7,537 academics in our datasets are provided in Table A1 in the Appendix.
We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table 2). The highest proportion of persistent users is observed within multidisciplinary researchers. The public metrics aggregated by field and by month are shown in Figure A1 in the Appendix. Researchers in the physical sciences and mathematics have notably higher average number of posts, indicating their consistent and significant contribution to discussions on Mastodon. Scholars in medicine and health sciences exhibit relatively less frequent posting.
Social & Behavioral Sciences 15.7 21.4 27.9 35.0 2918 38.72 98881 91958
Arts & Humanities 20.5 19.4 24.8 35.3 1522 20.19 38692 39986
Physical Sciences & Mathematics 12.0 21.2 27.8 39.0 851 11.29 11453 11980
Life Sciences 10.7 20.2 41.7 27.4 84 1.11 338 648
Education 21.3 26.2 24.2 28.3 244 3.24 2249 4171
Law 17.9 28.3 26.6 27.3 587 7.79 21664 27450
Medicine & Health Sciences 10.8 30.3 29.3 29.6 287 3.81 4292 4923
Multidisciplinary 10.8 18.3 28.8 42.1 1040 13.80 56371 52818
Business 25 25 25 25 4 0.05 60 66
Table 2. Distribution of user engagement and node degree statistics by academic field.
4.2. Mastodon Network Analyses
@@
To understand the connectivity within and between different fields, we collected the complete follower-followee data of all academic accounts on Mastodon, and constructed an internal connection network within each academic group, as well as a network aggregated by field of study.
4.2.1. Internal Network Aggregated by Field
@@
Figure 3 shows the network connections across different fields. In this network, node size represents the number of academics in each field, while edge width indicates the frequency of connections between the user accounts from different fields. Given the nature of the field, certain fields such as Social and Behavioral Sciences are strongly connected to Multi-disciplinary field. Our analysis reveals that the academic migration has led to the formation of a well-connected internal network, suggesting that researchers actively attempted to establish a new community on Mastodon. However, despite these coordinated efforts, the migration still faced significant challenges, as reflected through the low retention rates shown in Section 4.1.
4.2.2. Internal Networks aggregated by Server
Prior studies have shown that the Mastodon’s server instance based structure makes it easier for users to discover and connect with others within the same instance than outside instance. This leads to stronger within instance ties compared to weaker links across instances (Zignani et al., 2019). Therefore, we analyzed the follower-followee network aggregated by Mastodon instances, shown in Figure 4.
The academic users were spread across a total of 342 servers, with 9 server instances being isolates without any connecting edges to other instances, which were removed for clarity. Table A2 in the Appendix shows network centrality metrics for each server and the distribution of the users across different servers. It also includes the official description of the server instances, indicating whether they are customized to a specific group of scholars. During the migration, different disciplines used different strategies: some used general-purpose servers like mastodon.social which host a wide variety of content beyond just academic content, while others took a coordinated approach on setting up their field specific server instances such as infosec.exchange and journa.host. Figure 4 shows the follower-followee connections across different servers showing certain general purpose instances such as mastodon.social are popular among the tracked academic users (containing 21.3% of the total) than other general purpose servers like mas.to.

### v3

@@ Abstract
Following changes in Twitter’s ownership in 2022 and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some. Our study examines the behavior of a self-organized group of early academic adopters who joined Mastodon following changes in Twitter’s ownership. Utilizing publicly available user account data drawn from a voluntarily curated list of academics, we track the posting activity of these early adopters on Mastodon over a one year period. We also study follower-followee and interaction relationships to map internal networks, finding that the subset of academics who migrated to Mastodon were well-connected. However, this strong internal connectivity was insufficient to prevent users from returning to Twitter/X. Our analyses show that early adopters struggled to maintain engagement, shaped by Mastodon’s decentralized design and competition from alternatives such as Bluesky and Threads. The migration effort lost momentum after an initial surge, as most early adopters reduced activity or returned to Twitter. Our survival analysis further reveals that retention is strongly linked to diverse cross-server engagement and topic-server communities. Users with large pre-existing Twitter presence face significantly higher attrition risk, highlighting the challenge of replicating established social connections in decentralized ecosystem. By examining the coordinated migration attempt of early adopters, we find that even this highly motivated group faced substantial challenges, suggesting that later or less coordinated efforts would likely encounter even greater barriers.
Introduction
@@
Social media platforms have become essential outlets for communication, information dissemination, and professional networking. Among these platforms, Twitter [^1] Note on terminology: Twitter rebranded to X in July 2023, which occurred midway through our data collection. For simplicity, we use the terminology Twitter and tweets throughout the paper. (now known as “X”) has gained significant popularity and has been likened to a virtual town square for its role facilitating public discourse (Mascaro and Goggins 2012). Academics and journalists populate sub-communities on Twitter and use the platform for professional communication (Jurkowitz and Gottfried 2022; Malik et al. 2019).
In this work, we use the term Academic Twitter not to denote a formal organization, but as a shorthand for a diffuse network of scholars who use the platform for professional discourse. Academics have extensively used Twitter to share research findings, announce academic events such as conferences, and contribute to wider scholarly discourse (Veletsianos 2012; Mohammadi et al. 2018; Ebner 2013; Holmberg and Thelwall 2014). However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives (Cava et al. 2023).
@@
Among academics, Mastodon emerged as a popular alternative, with communities organized around shared interests. Unlike traditional platforms, Mastodon operates on a federated model in which independent servers communicate through a shared protocol, allowing users to join different instances while remaining connected across the network (Zignani et al. 2018). This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control (Raman et al. 2019). Yet despite these advantages, this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels (Raman et al. 2019).
@@
The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities. Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al. 2014). Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts of change (Fiesler and Dym 2020). Beyond platform-level triggers, prior studies highlight the role of social dynamics in shaping migration behavior. User migration has been linked to perceived platform attractiveness and peer effects (Hou and Shiau 2020). In the context of Twitter’s ownership change, migration to Mastodon reflected social influence, network structure, commitment, and shared identity, while remaining characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, and divergent posting behaviors across platforms (Cava et al. 2023; He et al. 2023). Recent work raises questions about sustainability, reporting lower engagement and posting activity on Mastodon compared to Twitter, and pointing to challenges in retaining active users within a decentralized architecture (Jeong et al. 2023).
@@
Despite prior work characterizing general trends across the population of platform users, there remains a lack of targeted research concerning migration of narrower and highly interconnected communities characterized by shared motivations and coordinated efforts. We address this gap by providing a targeted case study of early academic adopters responding to platform changes, examining their migration patterns and behaviors to offer a more nuanced understanding of how these users navigate platform transitions. Given the significant role that Twitter plays in many academics’ professional lives (Knight and Kaye 2016; Klar et al. 2020; Stewart 2015), academics represent a valuable subpopulation for studying migratory dynamics. Our work addresses the following research questions:
RQ1: How did early academic adopters navigate the migration to Mastodon, and what do longitudinal, network, and cross-platform patterns reveal about their retention?
RQ2: Which measurable user and community characteristics are associated with long-term retention on Mastodon?
@@
Our findings indicate that the majority of academic researchers who moved to Mastodon failed to maintain their posting activity and follower engagement over the long term. Although migrated academic accounts formed strongly-connected internal networks, these connections were insufficient to sustain user engagement over time. Of note, however, communities that established field specific server instances had higher user retention rate compared to those on general-purpose servers. Cross-platform analysis of most recent Twitter activity showed that most users returned to Twitter. Survival analysis confirms that user’s persistence is associated with a number of factors including diverse cross-server engagement, early platform activity, choice of server. Furthermore, cross-platform modeling shows a divergent pattern with long history of prior activity on Twitter is associated with higher attrition risks. These results underscore the challenges and opportunities associated with building and maintaining communities on new platforms.
Related Work
Cross-Platform Migration
Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al. 2014). Early work analyzed it using Push Pull Moore framework with Push effects (e.g., policy changes) drive users away, while Pull effects (e.g., better features) attract them to alternatives. However, Mooring factors such as switching costs often prevent complete migration (Chang et al. 2014; Lin and Wang 2017).
@@
Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process (Fiesler and Dym 2020). Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence (Hou and Shiau 2020). Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al. 2023). Analysis showed Twitter to Mastodon migration is characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al. 2023). A similar study on Twitter-to-Mastodon migration reveals lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al. 2023). These results are consistent with evidence that the transition is often an incomplete with users maintaining cross platform presence rather than fully substituting one platform to the other (Peña-Fernández et al. 2025). More broadly recent work on platform switching from Twitter shows that users disperse across multiple destination platforms rather than relocating to a single successor platform (Radivojevic et al. 2025). Our work looks more deeply at the specific case of the relatively well-synchronized and coordinated movement of communities of academic users across platforms.
Deplatforming as Involuntary Migration
Unlike voluntary platform migration, which is often motivated by user dissatisfaction or policy changes, deplatforming represents a form of forced migration where users are compelled to leave due to removal or suspension (Rogers 2020). A study of Reddit’s 2015 ban shows that deplatforming can reduce the visibility of harmful content on mainstream platforms while simultaneously displacing communities to alternative or less moderated spaces (Chandrasekharan et al. 2017). Parler was deplatformed after the 2021 U.S. Capitol attack, while Gab expanded as a refuge for banned Twitter users; however, research shows that such platforms often attract smaller or more fragmented audiences and face challenges related to moderation and community stability (Aliapoulios et al. 2021; Zannettou et al. 2018). While distinct from voluntary migration, deplatforming likewise underscores how platform-level conditions shape user mobility and contextualizes the broader challenges communities face when establishing presence on alternative platforms.
Characteristics of Mastodon
The decentralized structure of Mastodon and similar platforms leads to unique user engagement dynamics that are different from centralized platforms. Research suggests that users are drawn to Mastodon for the ease of protecting their information from data mining and the absence of algorithms shaping their feeds (Lee and Wang 2023). However, this can also make content discovery more difficult (Zignani et al. 2019) and may contribute to network fragmentation (Kleppmann et al. 2024), shaping interactions and communication patterns that differ from centralized platforms (Brauweiler et al. 2025). The decentralized structure encourages strong, tight-knit communities within individual server instances but fewer connections across instances (Zignani et al. 2018). These characteristics pose challenges for migration as well. Our study examines how specialized communities—particularly academics—navigated these dynamics in their movement from Twitter to Mastodon.
Data
@@
Our analyses are primarily based on a public Github repository containing lists of academics on Mastodon [^2] https://github.com/nathanlesage/academics-on-mastodon. These lists were curated by researchers from 50 distinct disciplines, who voluntarily added their own names and accounts. Each list corresponds to a research discipline, containing account details of researchers within that area, their Mastodon IDs, and a short description of their research. Some lists also included each user’s Twitter handle. For the purpose of longitudinal analysis, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022. That is, the users in our dataset were part of a coordinated, synchronized migration effort in November 2022.
@@
In total, we collected the Mastodon profiles and corresponding public metrics, e.g., number of followers, following, timestamps of most recent status of 7,542 academic user accounts using a customized Mastodon API [^3] https://github.com/halcy/Mastodon.py.git. We collected data weekly from November 2022 to October 2023 allowing for temporal analysis of user activity and engagement over the span of a year. In addition, we collected the most recent Mastodon data related to follower and following accounts for each scholar to conduct network analyses based on fields of study and associated servers. Additionally, to capture dynamics beyond follower and followee connections, we collected the user interactions (replies, boosts, mentions, and favorites) covering the full observation window from November 2022 to October 2023. We collected activity for users in our sample to extract both outgoing interactions and incoming interactions. Note that this interaction dataset was collected retrospectively in December 2025.
@@
Finally, to understand discussions surrounding Twitter migration to Mastodon on both platforms, we collected posts from both platforms during the one year window. Specifically, we collected tweets via Zeeschuimer [^4] https://github.com/digitalmethodsinitiative/zeeschuimer using the boolean search “mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01”. The set of academic-related terms was developed through a preliminary scan of tweets, ensuring coverage of both individual (e.g., “professor,” “scholar”) and collective (e.g., “academics,” “researchers”) identifiers. Since many tweets were included because the display name contained the word “mastodon” (as some users use their Mastodon account name as their Twitter display name) we further filtered out these tweets. We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets. We also collected posts through the Mastodon API using migration-related hashtags adapted from prior work (Jeong et al. 2024). We excluded the tag ”MastodonSocial” because it is commonly used for general content on Mastodon rather than migration. From a total pool of 12,654 migration related posts, we then applied a second stage filter using the same set of academic keywords described earlier to identify content relevant to academic migration. This process yielded 129 posts in the final dataset.
RQ1: Migration Dynamics and User Retention
@@
Following, we present findings from multiple analyses conducted to understand the dynamics of the migration. First, we used the longitudinal data from Mastodon to understand user attrition and engagement trends over the year. Then, we used network analysis to understand the connectivity and distribution of academics across Mastodon servers. Next, we conducted a cross-platform analysis using the subset of users who provided Twitter handles comparing their activity patterns on Twitter and Mastodon. Taken together, these analyses identify patterns and potential factors associated with user retention, which in turn motivate the selection of features included in the survival analysis for RQ2.
Longitudinal Patterns of Adoption and Retention
Attrition of academic accounts on Mastodon
@@
Figure 1 (Top) gives a bar plot illustrating the change in number of active users aggregated by month and a line plot showing the monthly retention rate based on the previous month. In this context, we considered a user to be active if they made at least a post in a given month. In the initial month of data collection (November 2022), there were 7,505 active scholars who had posted within that same month. [^5] There were 1, 4, and 32 scholars had their last activity in April, May, and October 2022, respectively, which fell outside our collection window. These scholars were excluded from the bar plot but included in subsequent analysis. Over the next seven months, the number of active scholars experienced a decline, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023. We observe a substantial rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform. The surge in July 2023 coincided with significant shifts in Twitter’s policy on limiting number of posts user can view and officially rebranded to “X” which may have prompted a renewed but temporary interest in Mastodon as alternative.
Engagement with academic accounts on Mastodon
@@
Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. Figure 1 (Bottom) presents the distribution of user active spans, which we use to characterize accounts into activity patterns for comparison across dimensions in later sections.
Figure 1: Top: Histogram showing number of monthly active users (active=at least one post per month). Bottom: Histogram showing distribution of user active spans (number of days they are on Mastodon aggregated by week).
@@
Active users. We define active users as scholars who engaged on Mastodon during the initial Twitter-to-Mastodon migration, which began in late October 2022. For our purposes, we’ve defined this initial period between October 15 and November 30, 2022. We categorize active users into four engagement groups based on the empirical distribution of active spans using the 25th, 50th, and 75th percentiles: one-time visitors; short-term, long-term, and persistent adopters. These data-driven cutoffs reflect the underlying heterogeneity in the duration of their activity on Mastodon following the migration. The days cutoffs are rounded to the nearest integer.
One-time visitors. These academics engaged only briefly at the beginning of the Twitter-to-Mastodon migration, with active spans falling below the 25th percentile of the overall distribution ($n_{days}\leq 39$). Their minimal participation suggests exploratory behavior during the migration wave without continued investment in the platform.
Short-term adopters. This group falls between the 25th and 50th percentiles of active spans ($39<n_{days}\leq 222$). These users engaged beyond initial exploration but did not sustain participation long enough to become stable contributors, potentially reflecting unmet expectations or competing commitments on other platforms.
Long-term adopters. These users fall between the 50th and 75th percentiles of the distribution ($222<n_{days}\leq 329$). They demonstrated extended engagement following the migration but did not remain active through the full analytic period.
@@
Persistent adopters. These users fall above the 75th percentile ($n_{days}>329$), corresponding to the longest active spans through the end of our analyses. Their sustained engagement suggests successful integration into Mastodon’s academic community and continued participation over time.
@@
Inactive users. Academics in this category were already using Mastodon before the migration wave from Twitter. Despite their early presence on the platform, they remained inactive during the migration period. One user was inactive during the first month but engaged briefly afterward. Inactive users are not included in subsequent analyses. The distribution of users is given in Table 1.
@@
Active One-time visitors 1895
Short-term adopters 1876
Long-term adopters 1921
Persistent adopters 1845
@@
Table 1: Number of users, by activity.
@@
We calculate the average following, follower and post counts for each user category, aggregated by month as shown in Figure A1. Our analysis reveals a sharp increase in all three public metrics at the onset of the migration period, specifically from November 2022 to December 2022. However, beyond this period, the rate of increase for following and follower metrics slows considerably. For one-time, short-term, and long-term users, growth in following and follower counts stagnated beginning in early 2023. As these users were not regularly contributing content, their visibility and potential to attract followers were likewise limited. In contrast, persistent users continued to post regularly. Despite their consistent activity, however, growth in their connections remained slow. This slow growth highlights the challenges faced in establishing and expanding social ties and communities on Mastodon. The persistent users’ continued efforts to engage and post suggest a commitment to the platform, yet the limited increase in followers indicates potential barriers in user retention and community building.
@@
Field-specific activity differences among academic accounts on Mastodon
@@
As discussed in Section Data, users self-identified their research disciplines in public lists. We aggregate these disciplines into 9 broad fields, consistent with existing taxonomies of academic disciplines (bepress. 2019). We adopt this taxonomy because its granular disciplinary structure aligns more closely with the self-reported research areas in our dataset, and it is widely deployed in institutional repository metadata system (e.g.,  [^6] https://guides.lib.odu.edu/c.php?g=502940&p=5839940, [^7] https://stars.library.ucf.edu/faq.html, [^8] https://digitalcommons.lib.uconn.edu/about.html). Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers. Inactive users were excluded from subsequent analyses. Disciplinary statistics of the 7,537 academics in our datasets are provided in Table A1 in the Appendix.
We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table A2 in the Appendix). The highest proportion of persistent users is observed among multidisciplinary researchers. One possible explanation is that interdisciplinary scholars engage with discussions spanning multiple fields, which may provide a broader set of conversational opportunities and connections that support continued participation. This observation motivates the inclusion of a binary indicator for interdisciplinary research in our subsequent modeling of user retention.
Network Structure and Community Formation
@@
To understand the connectivity within and between different fields, we collected the complete follower-followee network and multiple types of interaction networks of all academic accounts on Mastodon, and constructed an internal connection network within each academic group, as well as a network aggregated by field of study.
Internal networks aggregated by field
@@
Figure 2 shows the network connections across different fields. In this network, node size represents the number of academics in each field, while edge width indicates the frequency of connections between the user accounts from different fields. Given the nature of the field, certain fields such as Social and Behavioral Sciences are strongly connected to Multi-disciplinary field. Our analysis reveals that the academic migration has led to the formation of a well-connected internal network, suggesting that researchers actively attempted to establish a new community on Mastodon. However, despite these coordinated efforts, the migration still faced significant challenges, as reflected through the low retention rates shown in Section Longitudinal Analysis.